OpenStudy (anonymous):

solve the equation: log4(x-4)=-3 please show me step by step>.<

OpenStudy (anonymous):

4^(-3) = x - 4 x = 4^(-3) + 4

OpenStudy (anonymous):

just x=4^(-3)+4.. that's it???

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

simplified version of that would be x = 1/4^3 +4 x = 1/64 + 4 x = 1/64 + 4*64/64 x = 1/64 +256/64 x = 257/64